OpenStudy (anonymous):

what percent of 50 is 18

OpenStudy (bakonloverk):

\[\frac{ x }{ 100 }= \frac{ 18 }{ 50 }\] solve the proportion

OpenStudy (aakashsudhakar):

50 is half of 100. 100 represents a full percentage. This becomes much easier than it seems. 18 is to 50 as __ is to 100. Multiply 18 by 2 to get 36. Voila, it's 36%. 18 is 36% of 50.

OpenStudy (anonymous):

18 * 100 ---- 50

OpenStudy (anonymous):

\[\frac{ 50 }{ 18 }=\frac{ 100 }{ x }\] x=36